# AFBS:Buffer Gradient Selection in Semi-asynchronous Federated Learning

### v1

@@ Abstract
Asynchronous federated learning (AFL) accelerates training by eliminating the need to wait for stragglers, but its asynchronous nature introduces gradient staleness, where outdated gradients degrade performance. Existing solutions address this issue with gradient buffers, forming a semi-asynchronous framework. However, this approach struggles when buffers accumulate numerous stale gradients, as blindly aggregating all gradients can harm training. To address this, we propose AFBS (Asynchronous FL Buffer Selection), the first algorithm to perform gradient selection within buffers while ensuring privacy protection. Specifically, the client sends the random projection encrypted label distribution matrix before training, and the server performs client clustering based on it. During training, server scores and selects gradients within each cluster based on their informational value, discarding low-value gradients to enhance semi-asynchronous federated learning. Extensive experiments in highly heterogeneous system and data environments demonstrate AFBS’s superior performance compared to state-of-the-art methods. Notably, on the most challenging task, CIFAR-100, AFBS improves accuracy by up to 4.8% over the previous best algorithm and reduces the time to reach target accuracy by 75%.
@@
We are the first algorithm to perform gradient selection on semi-asynchronous federated learning buffers. By eliminating low-value stale gradients from the buffer, it significantly improves final accuracy and training speed. Moreover, the reduced number of gradients participating in aggregation decreases costly gradient summation operations, substantially lowering computational costs on the server side.
We propose a novel encryption algorithm. Initially, the client’s data distribution is transformed into a full-rank matrix through the addition of Gaussian noise, followed by information encryption via random projection. This method is also applicable to other scenarios requiring matrix encryption.
@@ 3.1 Preliminary
$\mathbf{v}_{k}^{(m)}=\mathbf{v}_{k-1}^{(m)}+\frac{1}{C}\sum_{i=1}^{C}\mathbf{v}_{i}^{(c)}\lambda.$ (7)
@@ 3.2 Motivation
The proposed method aims to mitigate the negative effects of outdated gradients in semi-asynchronous federated learning. We define the staleness of gradients as $\tau$, the difference between the current global round and the round in which the gradient source model was generated. An experiment shown in Figure 3 reveals that when a large number of outdated gradients are aggregated in each round, training performance significantly deteriorates, even worse than training with a single client. This highlights the need to remove outdated gradients from the buffer.
@@
Figure 4: The data distribution of pre-clustering versus non-pre-clustering strategies participating in training after gradient screening. In Fig.(a), we demonstrate the complete data distribution where each color represents a class. Fig.(b) shows the data distribution participating in training after pre-clustering, where the distribution patterns of each class remain largely unchanged. Fig.(c) presents the training data distribution without pre-clustering, revealing only 7 dominant clusters compared to the original 10 classes. This indicates that strategy (c) may lead to insufficient training for certain classes processed by slower servers.
@@ 3.5 Random projection encryption
This ensures that the distance between point pairs in high-dimensional space in the algorithm can still remain roughly unchanged in low dimensional space. Let us denote the label distribution matrix as $\mathbf{A}\in\mathbb{R}^{1\times d}$, and we perform the following operations to obtain a new matrix, where $\sigma$ is a specified tiny number.
@@ 4.1 Experimental Setup
We implement AFBS using PyTorch and FLGO [30], FLGO is a federated learning framework that supports AFL through time simulation. Our experiments are conducted on an NVIDIA GeForce RTX 4090 GPU.
@@
Datasets and Models. We test four datasets in total, covering both CV and NLP tasks. For CV, we test the LeNet5 model [31] on MNIST [32], CIFAR-10 [33], and CIFAR-100 [33]. For NLP, we test the FastText model [34] on SST2 [35]. We partition the dataset into K clusters using a Dirichlet distribution with $\alpha$ equal to 0.1. Within each cluster, the data distribution is independent and identically distributed, and the data volume follow a log-normal distribution with a standard deviation of 1.
@@ 4.2 Performance Comparison
We conduct experiments on all baseline methods under the condition of fixed virtual runtime and record the highest accuracy achieved by each algorithm, as shown in Table 2. It indicate that AFBS outperforms other methods significantly in terms of accuracy on the two most challenging datasets, CIFAR-10 and CIFAR-100. As shown in Figure 6, our algorithm consistently outperform the runner-up algorithm throughout the training process, with a leading margin of 4.26% and 4.8% respectively. Additionally, in the NLP task, the accuracy of AFBS is only slightly lower than FedDyn. This is because the number of clients deployed in the NLP task is only one-tenth of that in the CV task, making the impact of staleness relatively smaller in the NLP task. Nevertheless, even in this scenario, our method still outperform all asynchronous and semi-asynchronous algorithms in terms of accuracy. The above experimental results fully demonstrate that the proposed AFBS method exhibits outstanding performance under different levels of staleness impact.

### v2

@@ Abstract
Asynchronous federated learning (AFL) accelerates training by eliminating the need to wait for stragglers. However, its asynchronous nature introduces gradient staleness, where outdated gradients degrade performance. Existing solutions address this issue with gradient buffers, forming a semi-asynchronous framework. Nevertheless, this approach struggles when buffers accumulate numerous stale gradients, as blindly aggregating all gradients can harm training. To address this, we propose AFBS (Asynchronous FL Buffer Selection), the first algorithm to perform gradient selection within buffers while ensuring privacy protection. Specifically, the client sends the random projection encrypted label distribution matrix before training, and the server performs client clustering based on it. During training, server scores and selects gradients within each cluster based on their informational value, discarding low-value gradients to enhance semi-asynchronous federated learning. Extensive experiments in highly heterogeneous system and data environments demonstrate AFBS’s superior performance compared to state-of-the-art methods. Notably, on the most challenging task, CIFAR-100, AFBS improves accuracy by up to 4.8% over the previous best algorithm and reduces the time to reach target accuracy by 75%.
@@
We propose AFBS, the first algorithm to perform gradient selection on semi-asynchronous federated learning buffers. By eliminating low-value stale gradients from the buffer, it significantly improves final accuracy and training speed. Moreover, the reduced number of gradients participating in aggregation decreases costly gradient summation operations, substantially lowering computational costs on the server side.
We propose a novel encryption algorithm. Initially, the clients’ data distribution is transformed into a full-rank matrix through the addition of Gaussian noise, followed by information encryption via random projection. This method is also applicable to other scenarios requiring matrix encryption.
@@ 3.1 Preliminary
$\mathbf{v}_{k}^{(m)}=\mathbf{v}_{k-1}^{(m)}+\frac{1}{C}\sum_{i=1}^{C}\mathbf{v}_{i}^{(c)}\lambda_{i}.$ (7)
@@ 3.2 Motivation
The proposed method aims to mitigate the negative effects of outdated gradients in semi-asynchronous federated learning. We define the staleness of gradients as $\tau$, the difference between the current global round and the round in which the gradient source model was generated. An experiment shown in Figure 3 reveals that when a large number of outdated gradients are aggregated in each round, training performance significantly deteriorates, in some cases even falling below the performance of isolate single-client training. This highlights the need to remove outdated gradients from the buffer.
@@
Figure 4: The data distribution of pre-clustering versus non-pre-clustering strategies participating in training after gradient screening. In Fig.(a), we demonstrate the complete data distribution where each color represents a class. Fig.(b) shows the data distribution participating in training after pre-clustering, where the distribution patterns of each class remain largely unchanged. Fig.(c) presents the training data distribution without pre-clustering, revealing only 7 dominant clusters compared to the original 10 classes. This indicates that strategy (c) may lead to insufficient training for certain classes processed by slow clients.
@@ 3.5 Random projection encryption
This ensures that the distance between point pairs in high-dimensional space in the algorithm can still remain roughly unchanged in low-dimensional space. Let us denote the label distribution matrix as $\mathbf{A}\in\mathbb{R}^{1\times d}$, and we perform the following operations to obtain a new matrix, where $\sigma$ is a specified small constant.
@@ 4.1 Experimental Setup
We implement AFBS using PyTorch and FLGO [30], FLGO is a federated learning framework that supports AFL through time simulation. Our experiments are conducted on NVIDIA GeForce RTX 4090 GPU.
@@
Datasets and Models. We test four datasets in total, covering both CV and NLP tasks. For CV, we test the LeNet5 model [31] on MNIST [32], CIFAR-10 [33], and CIFAR-100 [33]. For NLP, we test the FastText model [34] on SST2 [35]. We partition the dataset into K clusters using a Dirichlet distribution with $\alpha$ equal to 0.1. Within each cluster, the data distribution is independent and identically distributed, and the data volume follows a log-normal distribution with a standard deviation of 1.
@@ 4.2 Performance Comparison
We conduct experiments on all baseline methods under the condition of fixed virtual runtime and record the highest accuracy achieved by each algorithm, as shown in Table 2. It indicates AFBS outperforms other methods significantly in terms of accuracy on the two most challenging datasets, CIFAR-10 and CIFAR-100. As shown in Figure 6, our algorithm consistently outperforms the runner-up algorithm throughout the training process, with a leading margin of 4.26% and 4.8% respectively. Additionally, in the NLP task, the accuracy of AFBS is only slightly lower than FedDyn. This is because the number of clients deployed in the NLP task is only one-tenth of that in the CV task, making the impact of staleness relatively smaller in the NLP task. Nevertheless, even in this scenario, our method still outperforms all asynchronous and semi-asynchronous algorithms in terms of accuracy. The above experimental results fully demonstrate that the proposed AFBS method exhibits outstanding performance under different levels of staleness impact.